Find rhe SNR of an uniform quantizer with 256 quantized level.
Answer #1

Answer :- The value of SNR is-

SNR = 6.02 *n + 1.76 dB

where n = log2(256) = 8, i.e. number of bits has been used.

Thus,

SŅR= 6.02 * 8 + 1.76 dB

SNR-49.92 dB